OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3). A. y = 3 B. y = 4x + 3 C. y = 4 D. y = 3x

OpenStudy (anonymous):

Substituting third equation into second equation, 3= 4m + (3-8m) 3= 4m +3 - 8m -4m= 0 m = 0 When m =0, from first equation, 3 = 8m +c 3= 8(0) + c c = 3 Therefore, y = mx +c y = (0)x + 3 y = 3 (Equation) so it was A Hope this Helps!